Find all complex numbers z such that z-=-32i, and give your answer in the form a+bi. Use the square root symbol V where needed to give an exact value for your answer. z = ???
Answer #1

Given,

z^{2}=-32i

Letz=a+bi

z^{2}=a^{2}-b^{2}+2abi=-32i

Therefore,

a^{2}-b^{2}=0.........(1) 2abi=-32i...............(2)

Also we have,

\left | z \right |^{2}=\left | z^{2} \right |

a^{2}+b^{2}=\left | -32i \right |

a^{2}+b^{2}=32....................(3)

Solving (1) and (3)

a^{2}=16

a=\pm 4

b=\pm 4

However, from (2) we come to know that both a and b cannot be negative at the same time

Therefore, (a,b) has 2 solutions (4,-4) and (-4,4)

Therefore solution is,

Z= 4-4i

or, Z= -4+4i